An aqueous solution contains 3.00% phenylalanine (C9H11NO2) by mass. Assume the phenylalanine is non-ionic and nonvolatile. What is the boiling point of the solution in degrees Celsius


1
Expert's answer
2021-11-06T01:30:06-0400

That means 100 g of solution contains 3 g of phenylalanine

Hence the mass of the water = 97 g

Mass of the phenylalanine = 3 g

Molar of the phenylalanine = 165 g/mol

The number of moles of phenylalanine "= \frac{3}{165} = 0.018 \;mol"

The molality of solution "= \frac{Number \; of \; moles \; of \;solute}{Mass \; of \; water \;(kg)}"

"= \frac{0.018\;mol}{0.097 \;kg} = 0.187 \;m"

The elevation in boiling point

"ΔT_{bp} = K_{bp} \times m \\

= 0.5121 \; °C/m \times 0.187 \;m \\

= 0.096 \;°C"

The boiling point of solution

"T_{bp}" = Boiling point of water + "\u0394T_{bp}"

"= 100 + 0.096 = 100.096 \; °C"